OpenStudy (luigi0210):

I need a little advice on this substitution problem:

OpenStudy (luigi0210):

So I have: \(\Large t^2 \frac{dy}{dt} +y^2=ty \) I got it on linear form for a Bernoulli sub: \(\Large \frac{dy}{dt} +\frac{y^2}{t^2}=\frac{y}{t} \) \( u=y^{1-2} =y^{-1} ; du=-y^{-2} dy \) Subbed in dy: \(\Large -y^2 \frac{du}{dt} +\frac{y^2}{t^2} =\frac{y}{t} \) Multiplied in u: \(\Large -y \frac{du}{dt} +\frac{y}{t^2} = \frac{1}{t} \) Got into linear again: \(\Large \frac{du}{dt}-\frac{1}{t^2} =-\frac{1}{t} y^{-1} \) Then from there it just gets weirder.. so I was wondering if I should have moved the \(ty\) to left and \(y^2 \) to the right in the beginning, what do you think?

ganeshie8 (ganeshie8):

try \(u = \frac{1}{y}\)

ganeshie8 (ganeshie8):

below is NOT standard form of bernouli - mistake : \[\frac{dy}{dt} +\frac{y^2}{t^2}=\frac{y}{t}\]

ganeshie8 (ganeshie8):

check again

OpenStudy (luigi0210):

It's \(\Large \frac{dy}{dx}+P(x)y=f(x) y^n \) right?

ganeshie8 (ganeshie8):

Yes

ganeshie8 (ganeshie8):

in your equation, power term and linear term got swapped

ganeshie8 (ganeshie8):

\[\frac{dy}{dt} -\frac{y}{t} = -\frac{y^2}{t^2}\] substitute \(u = \frac{1}{y^{2-1}}\)

OpenStudy (luigi0210):

So in short, yes, I should have switched, thanks gane :P

ganeshie8 (ganeshie8):

np :) see if you can watch 10 minutes of below video http://ocw.mit.edu/courses/mathematics/18-03-differential-equations-spring-2010/video-lectures/lecture-4-first-order-substitution-methods/ skip to 15th minute...

ganeshie8 (ganeshie8):

looks i have overlooked... your substitution is correct :)

ganeshie8 (ganeshie8):

after the substitution im getting this : \[\dfrac{d\color{Red}{u}}{dt} + \frac{\color{Red}{u}}{t} = \frac{1}{t^2}\] which is a linear eqn

ganeshie8 (ganeshie8):

which is exactly same as your last equation if you replace \(y^{-1}\) by \(u\)

OpenStudy (luigi0210):

I ended up getting \(\Large y^{-1}=\frac{ln ~t}{t} + c~t^{-1}\) Is the right? Answer key looks totally different P:

ganeshie8 (ganeshie8):

answers should match let me ask wolfram, one sec

OpenStudy (luigi0210):

The answer key gives \(\Large e^{\frac{t}{y}} =ct \)

ganeshie8 (ganeshie8):

both are right, lets see if we can convert your answer to the textbook form

ganeshie8 (ganeshie8):

\[y^{-1}=\frac{\ln ~t}{t} + c~t^{-1}\] which is same as \[\frac{1}{y}=\frac{\ln ~t}{t} +\frac{c}{t}\] multiplying through out by \(t\) \[\frac{t}{y}=\ln ~t +c\]

ganeshie8 (ganeshie8):

next take exponent both sides and define a new constant

ganeshie8 (ganeshie8):

\[\frac{t}{y}=\ln ~t +c\] takinng exponent both sides \[e^{\frac{t}{y}}=e^{\ln ~t +c}\]

ganeshie8 (ganeshie8):

\[e^{\frac{t}{y}}=e^{\ln ~t +c} = e^{\ln t}\cdot e^c = t\cdot e^c = tc_1\]

OpenStudy (luigi0210):

Yea, I realized how they did it after you flipped them around, thanks again, you're the best :D

ganeshie8 (ganeshie8):

<3